A project requires an initial investment of $4,000. The project is expected to generate positive cash flows of $2,500 a year for next three years and additional $300 in the last year (i.e., third year) of the project’s life. The required rate of return is 12%. What is the project’s net present value (NPV)? Based on the calculated NPV, should the project be accepted or rejected?
Answer #1

D7 fx =SUM(D3:06) 1 2 year Cash flows pv@12% Present value 3 0 $ (4,000) 1.0000 $ (4,000) 4 1 $ 2,500 0.8929 $ 2,232 2 $ 2,50

D7 fi =SUM(D3:06) с pv@12% A 1 2 year 30 4 1 5 2 Cash flows -4000 2500 2500 2800 =C3/1.12 =C4/1.12 =C5/1.12 NPV Present value